OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

In 1987 a recording artist sold 14,000 albums. In 1998 that same artist sold 4,683 albums. Find the rate of change in record sales from 1987 to 1998.

OpenStudy (anonymous):

A. 2003-09-02-00-00_files/i0300000.jpg years per album B. –847 albums per year C. 847 albums per year D. –847 years per album

OpenStudy (tkhunny):

Let's see YOUR work,. please. I'm sure Alex...bolt will be happy to do it for you, unless lessons have been learned, so why don't you beat A.b to it and show you YOUR work?

OpenStudy (anonymous):

what?

OpenStudy (tkhunny):

Simple enough. Show your work. Go!

OpenStudy (anonymous):

ummm

OpenStudy (alexandervonhumboldt2):

guees please not suspend me but he needs an answer and he already have th explanation 14000/-4683

OpenStudy (anonymous):

what is the answer?

OpenStudy (tkhunny):

If only that were the correct answer. Please try harder, both of you.

OpenStudy (alexandervonhumboldt2):

i dont undersand anything @dtuniyants

OpenStudy (anonymous):

uh I can't get it!

OpenStudy (anonymous):

me too. i don't know

OpenStudy (anonymous):

uh

OpenStudy (tkhunny):

\(\dfrac{Amount\;of\;Change}{Where\;We\;Started}\) In this case, we also passed 9 years, so divide by 9. You will first need the AMOUNT of the change.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so originally he sold 14000 is 1987 and 4683 in 1998. So change in sales = 14000 - 4683 = 9317 sales. rate of change = 9317 / 14000 x 100 %

OpenStudy (anonymous):

what is the answer @crazychickengirl123

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

@tkhunny is trying to explain, but you aren't participating.

OpenStudy (anonymous):

I don't know what tkhunny meant

OpenStudy (sleepyjess):

Then ask what they meant

OpenStudy (anonymous):

what does it mean?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

1,400,000

OpenStudy (anonymous):

that is not one of the answer choices

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

the answer is 1,400,000

OpenStudy (anonymous):

well all the answer choices are 847 and stuff

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

thats how many abuls he made

OpenStudy (mertsj):

\[\frac{4683-14000}{1998-1987}\]

OpenStudy (anonymous):

12010

OpenStudy (anonymous):

it wuled be a

OpenStudy (anonymous):

12010 @Mertsj

OpenStudy (anonymous):

a a a a a a aa a a a

OpenStudy (mertsj):

B

OpenStudy (anonymous):

thank you @Mertsj

OpenStudy (mertsj):

yw

OpenStudy (tkhunny):

@shifuyanli -- This is a collection of your contribution to this conversation. You show no work at all. You show no initiative. You show no cooperation. You show no background. You are simply begging for someone else to do your work for you. Eventually, someone did. You MUST do better than this. YOU must participate in obtaining the result or it is very, very unlikely to be of any real benefit to you or to anyone else. -- Tag -- Question -- Choices -- what? -- ummm -- What is the answer? -- uh I can't get it. -- uh -- okay -- what is the answer? -- Tag -- Tag -- Tag -- I don't know what tkhunny meant. -- What does it mean? -- Tag -- That is not one of the answer choices. -- Tag -- well all the answer choices are 847 and stuff -- Help! -- 12010 -- 12010 Tag -- Thank You Tag

OpenStudy (anonymous):

ik